OpenStudy (anonymous):

which one is the factor of 4u^4-16v^3 completely? A.4(u^2-2v)(u^2+2v) B.(2u-v)(2u+v)(u^2+v) C.(2u^2-4v)(2u^+4v) D.4(u^2-2v)(6^2-2v) i want to go with D but i think i am wrong

OpenStudy (anonymous):

Ok It must be C because I solved others and I am not getting answers from them

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

THe answer is A

OpenStudy (anonymous):

oops its not a

jhonyy9 (jhonyy9):

yes can be A if there is 16y^2 and not 16y^3

OpenStudy (anonymous):

its none of them unless you were looking for \[4(u^2-2v)(u^2+2v)=4(u^4+2uv-2uv-4v^2)=4u^4-16v^2\]

OpenStudy (anonymous):

oops should be \[+2u^2v-2u^2v\] in there

OpenStudy (anonymous):

It must be B sorry what I meant was B not C Because only B has three v

OpenStudy (anonymous):

Solve it You will get your answer or Maybe something wrong in the Question